90 lbs is 40% of a mix. How many lbs make up the other 60%? A 60/40 mix.
Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

90/0.4=225
225*0.6=135

Answer by stanbon(75887)

90 lbs is 40% of a mix. How many lbs make up the other 60%? A 60/40 mix.
------
Total mix = 90/0.4 = 225 lbs
----
Ans:: 0.6*225 = 135 lbs
--------------------
Cheers,
Stan H.
-------------